OpenStudy (steve816):

A container holds 67 g/mL. How many mg/pint is this?

jimthompson5910 (jim_thompson5910):

1000 mg = 1 g 1 pint = 473.176 mL (approx)

imqwerty (imqwerty):

1000 mg = 1 g 1 pint = 473.176 mL (approx) :)

OpenStudy (steve816):

Thanks nin, just what I needed!

OpenStudy (nincompoop):

\(\sf \huge 67 \frac{g}{mL} \times \frac{1000~mg}{1~g} \times \frac{473.176~mL}{1~pint} \)

OpenStudy (nincompoop):

how do you turn a rabbit into a frog \(\huge rabbit \times \frac{frog}{rabbit} \) how do you turn a frog into a human? \(\huge frog \times \frac{human}{frog} \) and that is why the princess kissed the frog and turned it into a prince

OpenStudy (nincompoop):

and then how do you turn the human back into a frog and then a rabbit? \(\huge \sf human \times \frac{frog}{human} \times \frac{rabbit}{frog} \)

imqwerty (imqwerty):

so the frog has to kiss the human now :)

imqwerty (imqwerty):

nd then the rabbit has to kiss the frog